### v1

# Beam-splitter-free, high-rate quantum key distribution inspired by intrinsic quantum mechanical spatial randomness of entangled photons

Abstract
Quantum key distribution (QKD) using entangled photon sources (EPS) is a cornerstone of secure communication. Despite rapid advances in QKD, conventional protocols still employ beam splitters (BSs) for passive random basis selection. However, BSs intrinsically suffer from photon loss, imperfect splitting ratios, and polarization dependence, limiting the key rate, increasing the quantum bit error rate (QBER), and constraining scalability, particularly over long distances. By contrast, EPSs based on spontaneous parametric down-conversion (SPDC) intrinsically exhibit quantum randomness in spatial and spectral degrees of freedom, offering a natural replacement for BSs in basis selection. Here, we demonstrate a proof-of-concept spatial-randomness-based QKD scheme in which the annular SPDC emission ring is divided into four sections, effectively generating two independent EPSs. pair photons from these sources, distributed to Alice and Bob, enable H/V and D/A measurements. The quantum-random pair generation inherently mimics the stochastic basis choice otherwise performed by a BS. Experimentally, our scheme achieves a 6.4-fold enhancement in sifted key rate, a consistently reduced QBER, and a near-ideal encoding balance between logical bits ‘0’ and ‘1.’ Furthermore, the need for four spatial channels can be avoided by employing wavelength demultiplexing to generate two EPSs at distinct wavelength pairs. Harnessing intrinsic spatial/spectral randomness thus enables robust, bias-free, high-rate, and low-QBER QKD, offering a scalable pathway for next-generation quantum networks.
keywords:
Entangled photon source, spatial-randomness, Bell’s parameter, density matrix, visibility, sifted key rate, QBERAyan Kumar Nai, Gopal Prasad Sahu, Rutuj Gharate, C. M. Chandrashekar G. K. Samanta
{affiliations}
Ayan Kumar Nai
Photonic Sciences Lab., Physical Research Laboratory, Ahmedabad 380009, Gujarat, India
Indian Institute of Technology Gandhinagar, Palaj, Gandhinagar 382055, Gujarat, India
Email Address: ayankrnai@gmail.com
Gopal Prasad Sahu
Photonic Sciences Lab., Physical Research Laboratory, Ahmedabad 380009, Gujarat, India
Indian Institute of Technology Dharwad, Dharwad 580011, Karnataka, India
Rutuj Gharate, G. K. Samanta
Photonic Sciences Lab., Physical Research Laboratory, Ahmedabad 380009, Gujarat, India
C. M. Chandrashekar
Quantum Optics Quantum Information, Department of Electronic and Systems Engineering, Indian Institute of Science, Bengaluru 560012, India
1 Introduction
Classical communication underpins the modern digital age, with data security relying on the computational complexity of breaking encryption schemes [1, 2]. However, the rapid progress in quantum computing, with scalable high-quality qubits on the horizon, poses a profound threat to this security. In particular, Shor’s algorithm [2, 3, 4] exposes the vulnerability of widely used classical encryption methods to quantum attacks, underscoring the urgent need for more resilient solutions.
Quantum key distribution (QKD), on the other hand, has emerged as a robust alternative for secure key exchange [5, 6, 7, 8]. QKD guarantees security even against attempts by an adversary to intercept the key [9, 10, 11], thereby remaining unaffected by advances in computational power, including those offered by quantum computers. As a result, the QKD protocols have progressed from laboratory demonstrations to real-world deployment through various protocols, including BB84 [1, 5], SARG04 [12], COW [13], and E91 [6]. Among these, BB84, the prepare and measure protocol based on weak coherent pulses, has emerged as the most practical choice for implementation, owing to its simplicity as well as the availability of composable security proofs [9, 14], despite known vulnerabilities to specific attacks [15, 16]. In parallel, entanglement-based QKD (EBQKD) protocols [6, 17, 18] have evolved, offering enhanced security owing to two fundamental quantum principles, the no-cloning theorem [19] and entanglement monogamy [20, 21, 22]. While EBQKD has been demonstrated over long free-space links by deploying entangled photon sources on satellites to enable key exchange between distant stations [8, 23, 24], its security, verified through entanglement measurements such as Bell inequality violations [6], ensures robustness against a wide spectrum of sophisticated attacks. Alternatively, the BBM92 protocol [25] enables secure key distribution without the need for continuous monitoring of Bell inequalities, offering higher key rates and reduced resource requirements provided that the shared state preserves maximal entanglement. In this framework, security is ensured through post-processed quantum bit error rate (QBER) evaluation combined with occasional Bell inequality verification, allowing device-independent operation [26] without relying on trusted nodes [17, 27]. This positions EBQKD as a strong candidate for future quantum networks. Nonetheless, practical implementations face inherent challenges, including low key rates originating from the limited brightness of spontaneous parametric down-conversion (SPDC) sources [28], and imperfections in 50:50 beam splitters used for random basis (Z and X) selection, thus introducing encoding bias and increased QBER [29]. Although brightness can be improved through the use of high figure-of-merit nonlinear materials and optimized system configurations [30, 31], the use of beam splitters remains a critical bottleneck. In particular, beam splitters not only suffer from practical deviations from the ideal 50:50 ratio, but also impose a fundamental 50 loss in conventional schemes [14, 32], thereby limiting achievable key rates. It is therefore crucial to explore new experimental strategies that achieve high key rates and low QBER while eliminating the dependence on beam splitters.
Here we demonstrate a high-performance implementation of the BBM92 protocol using a spatial-randomness-based encoding scheme. Polarization-entangled Bell states are generated from pair photons with annular emission in a type-0 phase-matched PPKTP crystal embedded in a polarization Sagnac interferometer. By exploiting the intrinsic quantum mechanical randomness of correlated pair photons located at diametrically opposite positions of the ring [33], our QKD approach obviates the need for a BS for random basis (Z, X) selection. Experimentally, we observe a 6.4-fold enhancement in the sifted key rate and a consistently reduced QBER compared with the conventional BBM92 protocol based on BS schemes. Moreover, the method achieves a near-ideal encoding balance between the Z (H/V) and X (D/A) bases, ensuring reliable key generation with low QBER.
2 Experiment
The concept of the spatial randomness–based BBM92 scheme is illustrated in Fig. 1(a). The entangled photon source (EPS) can generate any of the four Bell states, and . By applying a spatial-division scheme [34] to the annular ring distribution of the single EPS, constructed from a common set of physical resources including pump laser, nonlinear crystal, and interferometric architecture, we obtain two (see Fig. 1(b)) effective time-multiplexed EPSs (EPS 1 and EPS 2). The generation of pair photons from EPS 1 and EPS 2 is intrinsically quantum-random in time and unbiased. These pair photons are distributed between two users, Alice and Bob, with EPS 1 assigned to -basis measurements and EPS 2 to -basis measurements. Logical bits are encoded as the conventional schemes, with “0” and “1” corresponding to detections in the and polarizations, respectively. As a result, the intrinsic quantum randomness of photon-pair generation by EPS 1 and EPS 2 itself implements random basis selection, eliminating the need for an additional external beam splitter.
The schematic of the experimental setup is shown in Fig. 1(c). A continuous-wave, single-frequency diode laser delivering 40 mW of output power at a central wavelength of 405 nm, with a linewidth of approximately 20 MHz, is used as the pump source. A polarizing beam splitter (PBS) cube, in combination with a half-wave plate (HWP), is used to control the incident pump power. A second HWP, oriented with its fast axis at 22.5∘ relative to the vertical, rotates the horizontally polarized pump beam into a diagonal state. The pump is then focused to a waist diameter of 40 µm at the center of the nonlinear crystal using a plano-convex lens (L) with focal length mm. The nonlinear medium is a 10-mm-long periodically poled potassium titanyl phosphate (PPKTP) crystal (SLF Svenska Laserfabriken AB) with a mm2 in aperture and a single grating period m. The crystal, housed in a temperature-controlled oven (stability 0.1∘C, tunable up to 200∘C), is positioned at the center of a polarization Sagnac interferometer formed by a dual-wavelength PBS (D-PBS, designed for 405 and 810 nm) and two plane mirrors (M). A dual-wavelength half-wave plate (D-HWP) placed inside the Sagnac interferometer, with its fast axis at 45∘, flips the polarization between horizontal and vertical. Under these conditions, the crystal produces non-collinear, degenerate, type-0 () pair photons at 810 nm via quasi-phase-matched spontaneous parametric down-conversion (SPDC). As described in earlier work [31, 35], the clockwise and counterclockwise pump components inside the Sagnac interferometer generate pair photons with orthogonal polarizations, which combine at the D-PBS to produce the maximally entangled state, . The annular distribution of the down-converted photons is collimated by the lens (L) and separated from the pump by a dichroic mirror (DCM). The SPDC ring is divided into two halves using a prism-shaped gold-coated mirror (PM), and each half is further subdivided into two sections by D-shaped mirrors (DM), yielding four spatial regions denoted 1, 2, 1′, and 2′. For visualization, pair photons from regions (1, 1′) and (2, 2′) forming EPS 1 and EPS 2 are labeled as red triangles and green pentagons, respectively. The spatial sections (1, 2) and (1′, 2′) of the SPDC ring are distributed between Alice and Bob. The pair photons from EPS 1 (regions 1, 1′) are projected in the basis using PBS, while those from EPS 2 (regions 2, 2′) are projected in the basis after passing through HWP with its optic axis set at 22.5∘ relative to the vertical. The photons of the output ports of the PBSs are spectrally filtered by 3 nm interference filters (IF), coupled into single-mode fibers (SMFs), and measured by single-photon counting modules (SPCMs) connected to the time-to-digital converter (TDC) interfaced with a computer for coincidence detection and data processing. Both Alice and Bob have four SPCMs connected to the TDC to record the detection time stamps of photons in the four output ports corresponding to the mutually unbiased and bases. Unless otherwise stated, the coincidence window is maintained at 1 ns throughout the manuscript.
As observed previously [33], at any given instant EPS 1 and EPS 2 have equal probability of generating pair photons. Consequently, Alice and Bob measure photons either in the or in the basis, with negligible probability of simultaneous detections in both bases given the pump power used in the experiment. The encoding in the present scheme is illustrated in Fig. 1(d), in direct analogy to the conventional beam-splitter–based BBM92 protocol [29]. Here, the spatial randomness of photon-pair emission on the SPDC ring effectively replaces the role of the beam splitter, but without introducing bias. Like the standard BBM92 scheme, here also the detections in the basis due to the pair photon from EPS 1 are encoded as logical bits “0” and “1” for the and polarization outputs, respectively, and detections in the basis due to the pair photons of EPS 2 are encoded as “0” and “1” for the and outputs. As illustrated in Fig. 1(e), two entangled photon sources (EPS1 and EPS2) can be derived from a single source based on energy conservation, corresponding to the correlated signal–idler pairs and . The generation of pair photons with different wavelength combinations satisfying energy conservation is intrinsically random, same as the spatial distribution described above. Therefore, the conventional basis-selection element (BS) of the BBM92 protocol can be replaced with dichroic mirrors or wavelength-division multiplexing (WDM). Alice can encode information using photons at and for the Z and X bases, respectively, while Bob measures the corresponding partners at and for the Z and X bases. This approach enables the implementation of the BBM92 protocol at high bit rates without needing more than two channels for QKD between two users.
Finally, Alice and Bob announce their measurement bases over a public classical channel and retain only those events where the bases coincide, discarding the rest to generate the sifted key. The sifted key is subsequently processed using standard error correction (EC) and privacy amplification (PA) protocols to distill the final secure cryptographic key. It is important to emphasize that all experimental parameters reported here are derived directly from raw data without applying any corrections. While the implementation of such corrections could further improve the reported values, they do not hold practical significance for real-time applications of QKD systems. Hence, the uncorrected results presented here provide a more realistic performance assessment of the current protocol under operational conditions.
3 Results and discussions
We first characterized EPS 1 and EPS 2 in terms of their quantum performance parameters, including entanglement visibility, Bell parameter, and state fidelity. As both sources are derived from a single entangled photon source, they exhibit nearly identical characteristics. For clarity, we present the representative results of EPS 2 (see green paths in Fig. 1(b)). The measurements, summarized in Fig. 2, were performed at a fixed pump power of 3 mW and crystal temperature of T = 41.5∘C. We recorded the coincidence counts between photons transmitted through the PBS of both Alice and Bob, with Alice having fast axis of the HWP fixed at , , , and , corresponding to the (black), (blue), (red), and (green) bases, respectively, while Bob continuously rotated the HWP from to in steps. The resulting interference fringes are shown in Fig. 2(a).
As shown in Fig. 2(a), the coincidence counts exhibit the expected sinusoidal modulation, with measured fringe visibilities of 97.8, 97.8, 97.7, and 97.7 in the , , , and bases, respectively—well above the quantum threshold of 71 [31], thereby confirming high-quality entanglement. Furthermore, by analyzing coincidence counts across 16 different combinations of HWP settings of Alice and Bob and applying the standard CHSH inequality formalism [36], we obtained a Bell parameter of , clearly violating the classical bound by nearly 109 standard deviations. Further, we carried out quantum state tomography based on coincidence measurements across different polarization projections. The reconstructed real and imaginary parts of the two-qubit density matrix, shown in Fig. 2(b) and 2(c), respectively, confirm the generation of the maximally entangled Bell state with a fidelity of 97. The source brightness of EPS 2 is estimated to be 0.04 MHz/mW/nm. We observed comparable performance parameters for EPS 1. These results establish the suitability of both EPS 1 and EPS 2 for the implementation of the QKD protocol.
With the quantum parameters of the entangled photon sources established, we implemented the BBM92 protocol and benchmarked its performance against the conventional beam-splitter–based approach. Keeping the crystal temperature fixed at C, we measured the sifted key rate and QBER as functions of pump power. The results are shown in Fig. 3. As expected, the sifted key rate calculated by summing all the coincidence counts of same basis detection of two users (here, Alice and Bob), increases with pump power in both cases owing to the higher pair-generation rate. In our scheme (black dots), the sifted key rate rises from 0.29 Mbps to 2.9 Mbps as the pump power is increased from 1 to 13 mW, corresponding to a slope of 0.22 Mbps/mW for a one-meter free-space link. By contrast, the conventional scheme (black squares) exhibits only a modest increase from 0.045 Mbps to 0.47 Mbps, with a slope of 0.035 Mbps/mW. Such a superior performance of our current BBM92 scheme in terms of higher sifted key rate with respect to the conventional BBM92 scheme can be understood as follows. Let us consider one of the derived sources, EPS 1. In the absence of a beam splitter, the coincidence rate between Alice and Bob in the same basis (e.g., Z) is . Introducing an ideal 50:50 beam splitter at each photon path of Alice and Bob reduces the useful coincidences to the transmitted–transmitted and reflected–reflected outputs, leading to a 50% loss [37] and a key rate of . By contrast, our scheme avoids beam splitters and instead employs two sources, EPS 1 and EPS 2, yielding a combined rate of . Theoretically, this corresponds to a fourfold improvement, [M = (2)/(/2) = 4], in the sifted key rate compared with the conventional BBM92 protocol. Experimentally, we observe more than sixfold enhancement (Fig. 3(a)), which we attribute to non-ideal splitting ratios, inherent bias, and losses in practical beam splitters.
Further, we quantify the QBER using the following equations,
and
where and denote coincidence counts in the correct and incorrect bases, respectively. As shown in Fig. 3(b), the QBER in our scheme (black dots) increases from 1.03% to 6.44% at a slope of 0.45%/mW, whereas the conventional approach (black squares) rises from 1.8% to 8.04% at 0.55%/mW. These results confirm that the present BBM92 protocol, which exploits the intrinsic spatial randomness of SPDC photons, outperforms the conventional beam-splitter–based scheme by simultaneously achieving higher key rates and lower QBER.
We further verified the relation between entanglement visibility () and QBER () in our scheme, as this serves as a sensitive indicator of eavesdropping in the network. Previous works [39, 40] established that the Bell parameter and QBER are linked through , yielding the critical threshold of for secure key distribution [38]. Since also relates to the interference visibility as , the QBER and visibility are connected by . Experimentally, we varied the quantum interference visibility by tuning the pump power and measured the corresponding QBER using Eq. 1 and Eq. 2. The results are shown in Fig. 4. As expected, the QBER increases with decreasing interference visibility (averaged over four bases), with a slope of 1.78, close to the theoretical value of 2. This agreement confirms that visibility, and thus the derived Bell parameter, provides a reliable diagnostic for security in our current BBM92 protocol. Any tampering by an eavesdropper would reduce entanglement correlations, lowering below 2 and driving the QBER above the critical 11% threshold. Such conditions can be used to terminate the protocol before key generation, ensuring robustness of the system.
We compared the coincidence counts of Alice and Bob in the correct bases (, , , and ) for both our scheme and the conventional BBM92 protocol at a fixed pump power of 3 mW. The results are shown in Fig. 5. As evident from Fig. 5, in the conventional scheme, the coincidence rates are measured to be 3.2, 4.1, 3.1, and 2.7 kHz, respectively, while in our scheme the corresponding values increased to 20.3, 21.6, 21.6, and 21.8 kHz, respectively, showing an order-of-magnitude enhancement in the available bit rate. In addition to the higher count/bit rates, the distribution of coincidence events across the and bases provides a measure of the intrinsic randomness of the encoding. Ideally, the ratio should approach 1:1. For the conventional BBM92 protocol, we find a deviation to 1.3:1, reflecting imperfections of the beam splitter. In contrast, our scheme yields a nearly ideal ratio of 1:1, confirming that the spatially random distribution of SPDC photons ensures unbiased, quantum-mechanical encoding, as also shown in our previous report [34].
To provide a clear comparison between the present BBM92 scheme and the conventional beam-splitter–based BBM92 protocol, we summarize the key performance parameters—including QBER, sifted key rate, error-corrected key rate, privacy-amplified key rate, and the final secure key rate—in Table 1. The calculations employ standard post-processing procedures, with error correction implemented via low-density parity check (LDPC) codes and privacy amplification performed using established algorithms. As shown in Table 1, our spatial-randomness–based scheme consistently outperforms the conventional protocol across all stages of key generation, demonstrating its advantage for high-rate and secure quantum communication.
Parameters | Conventional | Present |
scheme | scheme | |
QBER (%) | 8.04 | 6.44 |
Sifted key rate (Mbps) | 0.47 | 2.90 |
Post EC key rate (Mbps) | 0.45 | 2.76 |
Key rate after PA (Mbps) | 0.15 | 1.02 |
Secure key rate (Mbps) | 0.15 | 1.02 |
4 Conclusions
We demonstrate a new experimental implementation of the efficient and scalable BBM92 QKD protocol by exploiting the quantum mechanical randomness of the spatial distribution of pair photons generated through momentum conservation in the down-conversion process. Using a polarization-entangled photon source with an annular-ring spatial distribution, we derived two entangled photon sources producing pair photons randomly and used for random basis selection, thereby replacing the beam splitter employed in conventional BBM92 schemes. Unlike the conventional approach, this method avoids bit loss due to the imperfections of the physical beam splitters. In a proof-of-concept demonstration over 1 m free space and under identical experimental conditions, the scheme achieved sifted key rate in megabits per second, a 6.4-fold enhancement, consistently lower QBER (7%) across varying pump powers, and a near-ideal encoding balance between logical bits ‘0’s and ‘1’s, thus eliminating encoding bias.
Although the current spatial demultiplexing scheme requires four quantum channels, it can be integrated into fiber-based QKD systems by exploiting energy conservation to generate two wavelength-correlated entangled sources from a single SPDC process and using wavelength demultiplexers in place of beam splitters. Thus, the spatial or wavelength demultiplexing-based randomness establishes a practical route toward high-rate, low-error, and bias-free entanglement-based BBM92 QKD protocol, paving the way for robust and scalable quantum communication networks.
ACKNOWLEDGMENTS
A. K. N., G. P. S, R.G., and G. K. S. acknowledge the support of the Department of Space, Govt. of India. A. K. N. acknowledges funding support for Chanakya - PhD fellowship from the National Mission on Interdisciplinary Cyber-Physical Systems of the Department of Science and Technology, Govt. of India, through the I-HUB Quantum Technology Foundation. G. K. S. acknowledges the support of the Department of Science and Technology, Govt of India, through the Technology Development Program (Project DST/TDT/TDP-03/2022).
AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
A. K. N. developed the experimental setup and performed measurements. G.P.S. participated in experiments, A. K. N., R.G., and C. M. C. participated in data analysis, numerical simulation, and data interpretation. A. K. N. and G. K. S. developed the ideas, and G. K. S led the project. All authors participated in the discussion and contributed to the manuscript writing.
DATA AVAILABILITY
The data that support the findings of this study are available from the corresponding author upon reasonable request.